7x+6y=-30
-8x+y=50
What is the solution of the system, using substitution method. I need an ordered pair for the answer, and I keep obtaining the wrong answer, somehow i got 2, 6 and that was not correct. Help me please! :)

Answer by nerdybill(7384)

Given:
7x+6y=-30 (equation 1)
-8x+y=50 (equation 2)
.
solve equation 2 for:
-8x+y=50
y=8x+50
.
substitute the above into equation 1 and solve for x:
7x+6y=-30
7x+6(8x+50)=-30
7x+48x+300=-30
55x+300=-30
55x=-330
x = -6
.
substitute above into 2 and solve for y:
-8x+y=50
-8(-6)+y=50
48+y=50
y = 2
.
answer: (x,y) = (-6, 2)